OpenStudy (anonymous):

question 6-The circuit elements have the following values: Vin=15V, R1=0.5Ω, R2=1.5Ω, R3=5Ω, A=2 and B=1.5!!!! anwers pls?????

OpenStudy (anonymous):

i1=Vo/(R3-B*R1*R3) Vo=Vin(R3-B*R1*R3)/(R3-(B*R1*R3)+R1+R2+AR2) Vth=Vo

OpenStudy (anonymous):

question 1 question 1- 1) (R*b)/(1+R*b)*vIN 2) (VS*(1+R*b)-vIN)/(1+R*b) current through R1 =current through R2 =(V1-R3(I1-I2))/(R1+R2+R3) current through R3 =(V1-(R1R2)(I1-I2))/(R1+R2+R3) 1. Va=(current through R2)*R2+Vb 2. Vb=(current through R3 )*R3 3. power in R1= current through R1*current through R1*R1 4. power in R2=current through R2*current through R2*R2 5. power in R3=current through R3*current through R2*R3 6. power in V1=V1*current through R1 7. power in i1=Vb*current through R3 8. power in i2=-Vb*current through R3 question 2 refer week 3 hw 3 q1 question 4 1. (R*b)/(1+R*b)*vIN 2. (VS*(1+R*b)-vIN)/(1+R*b) Question 5 Vgs=Vin 1) Vout=Vs-(f*(Vgs^3))*R 2) Gain = -3*f*(vgs^2)*R question 6 1)Vo* (R3-(B*R1*R3)+R1+R2+A*R2)/(R3-B*R1*R3)=VIN 2)i1=V0/(R3-B*R1*R3) 3)vo = vth 4)Rth=Vth/Isc